OpenStudy (anonymous):

omg someone help :( tell me how to solve it i wanna learn..thanks The are of a circle is 18(pi) square inches.What is the circumference of the circle to the nearest inch? use (pi)=3.14

OpenStudy (anonymous):

*area

OpenStudy (anonymous):

area = pi * r^2 circumference = 2 * pi * r r = radius find your radius first by using the area equation and then you'll be able to find the circumference. Does that help at all?

OpenStudy (anonymous):

ummm..ok so 3.14*r^2 but how would i find the radius with the area equation if i cant do anything else

OpenStudy (anonymous):

OHHH i think i got it

OpenStudy (anonymous):

you know that your area = 18 18 = 3.14 * r^2 so r is the square root of 18/3.14 ... so far so good?

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

i thought i had to multiply 18 by pi

OpenStudy (anonymous):

To get rid of pi on the right hand side you have to divide by 3.14 so you'll want do the same thing on the right hand side

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got 5.73

OpenStudy (anonymous):

so that's 18/pi ... now you need \[\sqrt{5.73}\] because right now it r^2 = 5.73 ... ok?

OpenStudy (anonymous):

ok i got 2.39

OpenStudy (anonymous):

great ... so now for your circumference 2 * pi * r (which is 2.39) and don't forget to add inches to your answer :)

OpenStudy (anonymous):

ok so i got 15.0092

OpenStudy (anonymous):

i can just put 15.92 right?

OpenStudy (anonymous):

You can't remove the zeros in between. They're important. :) Did your teacher tell you to round to any particular decimal place?

OpenStudy (anonymous):

oh ok, and no (: thank you so much

OpenStudy (anonymous):

Great ... I'd go with 15 inches myself. You are so very welcome! It was my pleasure.